OpenStudy (anonymous):

Find the exact value of sin -4pi/3 radians?

OpenStudy (anonymous):

so r u familiar with exact value graph

OpenStudy (anonymous):

\[\sin (2 \pi -\frac{ 4 \pi }{ 3 })\]

OpenStudy (anonymous):

\[\sin \frac{ 2 \pi }{ 3 }\]

OpenStudy (anonymous):

then you convert this into degree

OpenStudy (anonymous):

and you will get \[\sin 120^{0}\]

OpenStudy (anonymous):

the exact value then is \[\frac{ \sqrt{3} }{ 2 }\]

OpenStudy (welshfella):